OpenStudy (anonymous):

lim x->infinity (1-3/x)^(3x)

OpenStudy (anonymous):

Any help? i know the answer just stuck on a certain part.

OpenStudy (anonymous):

y = lim (1-3/x)^3x log both sides ln y = lim ln (1-3/x)^3x ln y = lim 3x * ln(1-3/x)

OpenStudy (anonymous):

Here are the terms I expanded to. Then you can see the pattern...

OpenStudy (anonymous):

use l'hopital's rule, first you must rewrite to infinity / inifity form ln y = lim ln(1-3/x) / (1 / 3x)

OpenStudy (anonymous):

differentiate top and bottom, it simplifies to ln y = lim -9 / (1 - 3/x) now the limit is obviously -9 ln y = -9 y = e^-9 = the original limit

OpenStudy (anonymous):

This is what i've gotten so far, lim x-> infinity (1-3/x)^(3x) y= (1-3/x)^(3x) ln y = (3x) * ln(1-3/x) lim x->infinity 3x * ln (1-3/x) I have no idea how to do it.

OpenStudy (anonymous):

with l'hopital's rule

OpenStudy (anonymous):

but 3x * ln(1-3/x) is in the form Infinity * 0 so you can't do it just yet. you can move the 3x into the denominator

OpenStudy (anonymous):

so pretty much lim x-> infinity (1-3/x)/(1/3x) then do i use L'hopitals rule?

OpenStudy (anonymous):